### v1

@@ II-B1 Applications of LLM Agents
LLMAs are used for different applications. One application that has especially gained traction is programming, primarily because of LLMs ability to generate code given a text prompt. LLMAs are used to automate different parts of the software lifecycle, including development, deployment, testing, and fixing bugs [49]. Other applications propose using LLMAs in healthcare for counselling [28], financial trading [47], automating customer service [66] and shopping assistants [81].
@@ II-B2 Models for LLM Agents
The different components used in the standard model of a LLMA include a memory, retrieval mechanism, action sets, and an environment.  [71] studies different cognitive architectures using these components for LLMAs. And there has been a lot of work to improve the capabilities of these components in LLMAs, using a dynamic context [15] and using self notes to perform continual learning [45]. However, we propose augmenting the LLM with a Bayesian engine to perform sequential Bayesian inference on a stream of data. The Bayesian engine model proposed in this paper can also be used for more general tasks, as we discuss in the conclusion.
@@ II-B3 Networks of Agents
Since many of the LLM agents interact with other agents directly or through content generated by them, there is a need for more systematic and mathematically rich black-box models for LLMs and LLM agents (LLMAs). Such models help understand their behavior and eventually control it to ensure reliability. Often, the collaboration of LLMAs are modeled as a graph or network [24], which  [52] proposes dynamically adapting depending on the task. There are various different programmatic frameworks where the LLMAs can be abstractly programmed to perform different tasks [77, 80]. Some of these frameworks even allow making these agents autonomous [86]. The methods in this paper deal with a line graph topology of LLMAs, which perform sequential Bayesian estimation. The setup studied in this paper can be extended to more general graph topologies, and different issues, including data-incest, can be studied.
@@ II-C1 Multiple Bayesian agents sequentially estimating a state
The motivation for multiple such Bayesian agents, each receiving private observations, is motivated by privacy, improved detection, and finite context length. If the same private observation (even the low-dimensional representation) is used, the LLM can be fine-trained on this data, which might contain sensitive information [20]. Also, different LLMs can be given a diverse set of contexts, which enables reducing the bias involved with their decisions [35]. Also, practically due to the finite context length, the observations can be considered private with respect to consecutive LLMAs evaluations.
@@ II-C2 Framework of Bayesian social learning
Recently [56] looked at social learning in LLMs using a teacher-student framework, but this work was in a static setting where the LLMs don’t have a belief that they adaptively update. In general, sequential social learning in Bayesian agents has been studied extensively [13], and our work formalizes the problem of Bayesian social learning in LLMAs. The theoretical results presented in this paper have been studied before in the context of distributed Bayesian sensors in [39] and [7]. Compared to [7, 39], we look at the LLMAs as interpretable Bayesian sensors and provide a more comprehensive outlook. We also look into multiple observations being shared and provide a concentration inequality for overspending the incentive with respect to a budget constraint. Recently [46], looked at detecting information cascades using deep learning. Although in this paper we focus on delaying an information cascade to improve estimation accuracy, methods similar to [46] can be integrated with our approach.
@@ II-D Interpretability and Social Fairness
However, we use an interpretable Bayesian model to perform sequential Bayesian inference from text observations of an underlying state, whereas [69] create an xAI model with decision trees and n-grams models using outputs from an LLM. Another such work is [72], where the authors propose training a separate concept neural network that uses the output of an LLM to interpretably classify text embeddings. This approach can complement the work in this paper when the setting is dynamic. Our work uses tools from revealed preferences and social learning to analyze the behavior of individual and interacting LLMAs from a microeconomic lens.
Further, our focus is also on building LLMAs, which are safe, reliable, and fair, goals which are aligned with operationalizing responsible AI [87]. This becomes challenging to do since LLMAs have been known to show biases which are inherent to human beings like conformity [6] and bias towards different attributes [9]. These effects will be more prominent when the LLMAs interact with each other in different scenarios like a Mindstorm [88] or a language model cascades [25]. Finally, the need to study interpretable models for LLMAs is motivated by the rise of unified agents [59], which are a representation of a trend in artificial intelligence that the different models are converging to a single efficient model [29]. This paper therefore tries to systematically understand Bayesian social learning in LLMAs, to help prevent undesirable phenomena like model collapse [67].
There has been substantial work in the fairness of the machine learning models [53], and even evaluating large language models for different measures of social fairness [14]. However, [3] recently highlighted how it is extremely difficult to benchmark LLMs on existing fairness metrics because of the way LLMs are used. This becomes even more challenging for LLMAs, where the agents further have agency and can also communicate with other LLMAs. Therefore, the focus of our study is to construct interpretable models for LLMAs, which can be used to understand the decisions of LLMAs. This understanding can help construct more suitable societal fairness metrics. Our work additionally has relatively mild assumptions on the utility/cost function of the LLMAs, and hence can be adapted for different sociological costs.
@@ II-E3 Product Quality Identification
One of the issues on e-commerce platforms is to identify bad products early on; however, just using numerical ratings can be uninformative, especially when the number of ratings is less. However, there is a lot of information contained in the descriptive reviews of the product, which one can efficiently extract using LLMs. Therefore, a sequence of LLMAs can efficiently analyze the reviews of a product to identify if the product is of good quality or not. This is an extended and sophisticated version of opinion mining, which has shown to be effective in sentiment analysis on online platforms [50].
@@ III LLM Agent as a Social Sensor
The different aspects of the mathematical model for the LLMAs are discussed, and the utility of a LLMA is introduced, which can be reconstructed for a black-box LLMA using the Bayesian revealed preference framework discussed later. We discuss how the framework and the results of the paper can be extended to contemporary models like vision language models (VLMs). In essence, this section, therefore, shows how LLMA acts as a social sensor, which can be applied to sophisticated settings such as online platforms where physical sensors don’t work to sense the underlying state from observation obtained by the interaction of humans.
@@
We further motivate the construction of such a model from the point of view of interpretability, reliability, and controllability. There has been a lot of work done to improve the interpretability of the output of a standard LLM [70]. In the black-box setting, the approach proposes asking the LLMs to provide a reason in addition to the output [84]. This works well in practice for simple applications, however when sequential Bayesian inference needs to be performed on millions of text observations interpreting the reason itself becomes a tedious task. Therefore, we propose using the LLMs as a low-dimensional map from the high-dimensional text space by designing prompts that are useful in analyzing. The LLMs can either be explicitly controlled using the system prompt or their outputs can be restricted to a certain state.
A Bayesian engine is then used to provide confidence on these low-dimensional variates, which is an easier task to do than on the high-dimensional text data due to the curse of dimensionality. This helps in using the LLMA in a reliable way since the LLMAs can provide confidence in the actions they would take given the observations. Such a model of LLMA is also controllable with respect to the cost function associated with the Bayesian engine, as we illustrate below.
Notation: Boldface symbols (e.g., $\bf{x}$) denote a row vector, and $\bf{x}^{\prime}$ denotes the transpose of the vector. $\mathsf{diag}(\bf{x}^{\prime})$ denotes a diagonal matrix with $\bf{x}$ as diagonal entries. Capital letters (e.g. $B$) denote matrices and $B_{y}$ denotes $y^{\text{th}}$ row of the matrix.
III-B Mathematical Model for a Utility Maximizing LLMA Performing Bayesian Inference
We consider a large language model agent composed of a large language model (LLM) and a Bayesian Engine.
III-B1 Mathematical Model for LLM as sensor
First, we give a mathematical model for a general-purpose LLM. For this paper, we consider black-box access to the LLM and hence both open-source LLMs like Llama3 [74] and closed-source LLMs like ChatGPT. One of the ways to model a black-box LLM is as an input-output block. The input to an LLM is a single text prompt, which we decompose into three things: the system prompt, which we refer to as control, the context, and the user prompt, which we refer to as observation.
Assume that the dictionary of all words of the LLM (tokens  [^4] Although the input of the LLM is text, in most of the architectures, the text (or more generally multi-modal data) input is first decomposed into different tokens and the tokens are processed, but since we consider the LLM as a black-box, our interpretable models abstract these implementation details.) is given by $\mathcal{D}$, and this dictionary also includes the blank word. The time index is given by $k=1,2,\dots$. A black-box LLM can be viewed as an input-output block.
@@ III-B1 Mathematical Model for LLM as sensor
Following the system prompt, the next input is a context to the LLM. This context could contain external information and examples that is time-dependent and may depend on previous interactions, which is dynamic and can not be put as a part of the system prompt. We consider a context of length $m_{\text{context}}$ and at time $k$ denote it by $\kappa_{k}\in\mathcal{D}^{m_{\text{context}}}$.
@@
We consider an LLM, which is pre-trained on trillions of tokens of text to autoregressively generate the next-token  [^5] There are other techniques of generating the token, but they can be accommodated by a suitable augmentation in the mathematical formulation presented here.. For developing a token of length 1, the output of the LLM can be given by a conditional probability distribution $\phi(c_{k},\kappa_{k},z_{k})=\mathds{P}(\cdot|c_{k},\kappa_{k},z_{k})$, where $z\in\mathcal{Y}^{\prime}$ is the user prompt, $c_{k}$ is the system prompt, $\kappa_{k}$ is the context, and the function $\phi:\mathcal{D}^{m_{\text{control}}}\times\mathcal{D}^{m_{\text{context}}}\times\mathcal{D}^{m_{\text{user}}}\to\mathcal{P}(\mathcal{D})$ outputs a probability distribution over the dictionary $\mathcal{D}$. For generating an output of an LLM with length $>1$, we consider a function $g$ which takes in the function $\phi$ and outputs tokens from the space $\mathcal{D}^{m_{\text{output}}}$, where $m_{\text{output}}$ is the maximum length of the output. The output of the LLM, denoted by $y$ is obtained as $y=g(\phi,c_{k},\kappa_{k},z_{k})$. Therefore, we can represent a black box LLM with the following tuple,
@@
    The LLMs in production today don’t explicitly provide confidence in their estimates. Confidence is important, especially when deploying these LLMs in situations where the safety and reliability of the system are important, e.g., financial and healthcare applications.
@@
    The confidence in the estimates is also useful when these LLMs are deployed in human-in-the-loop systems, where a human expert can be called for in case the confidence that the LLMs have in their estimates is not high enough. Relying solely on a single LLM is not reliable since the LLMs are fallible to adversarial prompting.
@@
To this end, we propose using the LLM as a map from the high-dimensional space $\mathcal{Y}^{\prime}$ to a low-dimensional space $\mathcal{Y}$ which takes value from whose cardinality $Y\ll\mathcal{D}^{m_{\text{user}}}$. We construct a Bayesian engine composed of two probability distributions, the prior of the state, $\pi\in\mathcal{P}(\mathcal{X})$ and the likelihood of a low-dimensional observation $D:\mathcal{X}\to\mathcal{P}(\mathcal{Y})$ and is a conditional probability distribution. We use $B$ and $D$ to denote the observation matrices in the rest of the paper.
@@
We provide the following examples illustrating how LLMs can be used as maps from the text space to a low-dimensional space. We discuss how this low-dimensional space can be constructed by an analyst or engineer so that the posterior computed by the Bayesian engine is interpretable.
@@ III-B2 Construction of Bayesian Engine
Figure 5: LLMAs can be used to interpret a financial indicator (difference of close prices) by parsing financial news to extract 1616 1616 interpretable features in Example 2. We analyze the news articles from $03/2020$ to $08/2020$ corresponding to the AAPL stock. We query the LLM for 1616 1616 binary with different features, including whether the news article indicates optimism about the market and whether there is investor interest in the stock. The interpretable features can be used for analysis of the stock (and subsequently for Bayesian inference), as illustrated by the difference in the ratio of the stock prices across days.
@@
For a state $x\in\mathcal{X}$, the LLMA receives a text observation $y^{{}^{\prime}}\in\mathcal{Y}^{\prime}$ which the LLM of the LLMA parses and provides a low-dimensional output $y\in\mathcal{Y}$. The LLMA has a Bayesian engine which has a prior on the state space given by $\pi$ and an observation likelihood $D$. The LLMA uses the low-dimensional output from the LLM to compute the posterior using Bayes’ rule,
$\displaystyle\mathds{P}_{D,\pi}(x|y)=\frac{D_{y,x}\pi(x)}{\sum_{x}D_{y,x}\pi(x)}.$
@@
$\displaystyle u=\underset{{u\in\mathcal{U}}}{\arg\max}\sum_{x\in\mathcal{X}}r(x,u)\mathds{P}_{D,\pi}(x|y).$
@@
$\displaystyle LLMA(\mathcal{L},\pi,D,r),$ (2)
where $\mathcal{L}$ is an LLM of the form (1), $\pi$ is the prior over the state, $D$ We make the following remarks on our interpretable sensor model construction of a LLMA.
@@
We described the operation of Bayesian inference on a single observation when there is a stream of observations $y_{1},\dots,y_{k}$ then the prior is updated after every step with the computed posterior $\pi_{k+1}=\mathds{P}_{D,\pi_{k}}$.
@@
The likelihood $D$ can be computed by using the LLM on a set of synthetic or public offline data where we simulate the state and use the text observations to obtain the low-dimensional observations from the LLM.

### v2

@@ II-B1 Applications of LLM Agents
An important application LLMAs are used for is programming, primarily because of LLMs ability to generate code given a text prompt. LLMAs are used to automate different parts of the software lifecycle, including development, deployment, testing, and fixing bugs [26]. Other applications propose using LLMAs in healthcare for counselling [28], financial trading [27], automating customer service [17] and shopping assistants [31].
@@ II-B2 Models for LLM Agents
The different components used in the standard model of a LLMA include a memory, retrieval mechanism, action sets, and an environment.  [32] studies different cognitive architectures using these components for LLMAs. There has been a lot of work to improve the capabilities of these components in LLMAs, using a dynamic context [42] and using self notes to perform continual learning [61]. However, we propose augmenting the LLM with a Bayesian engine to perform sequential Bayesian inference on a stream of data. The Bayesian engine model proposed in this paper can also be used for more general tasks, as we discuss in the conclusion.
@@ II-B3 Networks of Agents
Since many of the LLM agents interact with other agents directly or through content generated by them, there is a need for more systematic and mathematically rich blackbox models for LLMs and LLM agents (LLMAs). Such models help understand their behavior and eventually control it to ensure reliability. Often, the collaboration of LLMAs are modeled as a directed graph [25, 36, 37], which  [34] proposes dynamically adapting depending on the task. There are various different programmatic frameworks where the LLMAs can be abstractly programmed to perform different tasks [62, 35]. Some of these frameworks even allow making these agents autonomous [14]. The methods in this paper deal with a line graph topology of LLMAs, which perform sequential Bayesian estimation. The setup studied in this paper can be extended to more general graph structures, and different issues such as data incest can be studied.
@@ II-C1 Multiple Bayesian agents sequentially estimating a state
The interaction of multiple such Bayesian agents, each receiving a private observation, is motivated by privacy, improved detection, and finite context length. If the same private observation (even the low-dimensional representation) is used, the LLM can be fine-trained on this data, which might contain sensitive information [39]. Also, different LLMs can be given a diverse set of contexts, which enables reducing the bias involved with their decisions [19]. Also, practically due to the finite context length, the observations can be considered private with respect to consecutive LLMAs evaluations.
@@ II-C2 Framework of Bayesian social learning
Recent work has looked at social learning in LLMs using a teacher-student framework [40], but this work was in a static setting where the LLMs do not have a belief that they adaptively update. In general, sequential social learning in Bayesian agents has been studied extensively [3], and our work formalizes the problem of Bayesian social learning in LLMAs. The theoretical results presented in this paper have been studied before in the context of distributed Bayesian sensors in [4] and [2]. Compared to [2, 4], we view the LLMAs as interpretable Bayesian sensors and provide a more comprehensive outlook. We also consider multiple observations being shared and provide a concentration inequality (Theorem 5) for overspending the incentive with respect to a budget constraint. Recently [63], looked at detecting information cascades using deep learning. Although in this paper we focus on delaying an information cascade to improve estimation accuracy, methods similar to [63] can be integrated with our approach.
@@ II-D Interpretability and Social Fairness
However, we use an interpretable Bayesian model to perform sequential Bayesian inference from text observations of an underlying state, whereas [45] create an xAI model with decision trees and n-grams models using outputs from an LLM. Another such work is [65], where the authors propose training a separate concept neural network that uses the output of an LLM to interpretably classify text embeddings. This approach can complement the work in this paper when the setting is dynamic. Our work uses tools from revealed preferences and social learning to analyze the behavior of individuals and interacting LLMAs from a microeconomic lens.
Further, our focus is also on designing LLMAs, which are safe, reliable, and fair, goals which are aligned with operationalizing responsible AI [51]. This becomes challenging to do since LLMAs have been known to show biases which are inherent to human beings like conformity [66] and bias towards different attributes [67]. These effects will be more prominent when the LLMAs interact with each other in different scenarios like a Mindstorm [33] or a language model cascades [68]. Finally, the need to study interpretable models for LLMAs is motivated by the rise of unified agents [69], which are a representation of a trend in artificial intelligence that the different models are converging to a single efficient model [70]. This paper therefore tries to systematically understand Bayesian social learning in LLMAs, to help prevent undesirable phenomena like model collapse [15, 38].
There has been substantial work in the fairness of the machine learning models [48], and even evaluating large language models for different measures of social fairness [71]. However, [49] recently highlighted how it is extremely difficult to benchmark LLMs on existing fairness metrics because of the way LLMs are used. This becomes even more challenging for LLMAs, where the agents further have functionality and can also communicate with other LLMAs. Therefore, the focus of our study is to construct interpretable models for LLMAs, which can be used to understand the decisions of LLMAs. This understanding can help construct more suitable societal fairness metrics. Our work additionally has relatively mild assumptions on the utility/cost function of the LLMAs, and hence can be adapted for different sociological costs.
@@ II-E3 Product Quality Identification
One of the issues on e-commerce platforms is to identify poor quality products early on; however, just using numerical ratings can be uninformative, especially when the number of ratings is less. However, there is a lot of information contained in the descriptive reviews of the product, which one can efficiently extract using LLMs. Therefore, a sequence of LLMAs can efficiently analyze the reviews of a product to identify if the product is of good quality or not. This is an extended and sophisticated version of opinion mining, which has shown to be effective in sentiment analysis on online platforms [53].
@@ II-E4 Personalized Recommendation Systems
II-F Comparision To Prior Work
The primary focus of this paper is to employ ideas from statistical signal processing and microeconomic theory to model and analyze interacting LLMAs. However, given the domain-specific constraints that are inherent to LLMAs, we claim the following modeling and theoretical innovation over prior work in Bayesian social learning and interpretable utility reconstruction: a) There is a hierarchical observation structure inherent to the Bayesian sensor model of an LLMA considered here, whereas the Bayesian revealed preference framework only has a single observation likelihood [1] b) There has been limited research on interpretable models for LLMAs. However, in contrast to the other interpretable machine learning literature, including glassbox models, we reconstruct a utility function and an information acquisition cost given only blackbox access to the LLMA and is more closely aligned to [46]. c) We study the effect of sharing more private observations and a better observation likelihood on herding in a classical Bayesian social learning setting [3, 75] d) The social learning framework considered here allows for LLMAs with different LLMs and is presented for the specific application of Bayesian inference in contrast to [40].
@@ III LLM Agent as a Social Sensor
The different aspects of the mathematical model for the LLMAs are discussed, and the utility of a LLMA is introduced, which can be reconstructed for a blackbox LLMA using the Bayesian revealed preference framework discussed later. We discuss how the framework and the results of the paper can be extended to contemporary models like vision language models (VLMs). In essence, this section, therefore, shows how LLMA acts as a social sensor, which can be applied to sophisticated settings such as online platforms where physical sensors do not work to sense the underlying state from observation obtained by the interaction of humans.
@@
We further motivate the construction of such a Bayesian sensor model for an LLM from the point of view of interpretability, reliability, and controllability. There has been a lot of work done to improve the interpretability of the output of a standard LLM [59]. In the blackbox setting, the approach proposes asking the LLMs to provide a reason in addition to the output [60]. This works well in practice for simple applications, however when sequential Bayesian inference needs to be performed on millions of text observations interpreting the reason itself becomes a tedious task. Therefore, we propose using the LLMs as a low-dimensional map from the high-dimensional text space by designing prompts that are useful in analyzing. The LLMs can either be explicitly controlled using the system prompt or their outputs can be restricted to a certain state.
A Bayesian engine then uses these low-dimensional variates to update the belief, which is an easier task to do than on the high-dimensional text data due to the curse of dimensionality. This helps in using the LLMA in a reliable way since the LLMAs can provide confidence in the actions they would take given the observations. Such a model of LLMA is also controllable with respect to the cost function associated with the Bayesian engine, as we illustrate below.
Notation: ${x}^{\prime}$ denotes the transpose of vector $x$. $\mathsf{diag}(\bf{x}^{\prime})$ denotes a diagonal matrix with $\bf{x}$ as diagonal entries. Capital letters (e.g. $B$) denote matrices and $B_{y}$ denotes $y^{\text{th}}$ row of the matrix.
III-B Bayesian Utility Maximization Model for LLMA
We consider a LLMA composed of a large language model (LLM) and a Bayesian Engine.
III-B1 Abstracting a Large Language Model (LLM) as a sensor.
First, we give a mathematical model for a general-purpose LLM. For this paper, we consider blackbox access to the LLM and hence both open-source LLMs like LLaMA3 [58] and closed-source LLMs like ChatGPT. One of the ways to model a blackbox LLM is as an input-output block. The input to an LLM is a single text prompt, which we decompose into three things: the system prompt, which we refer to as control, the context, and the user prompt, which we refer to as observation.
Assume that the dictionary of all words of the LLM (tokens [^5] Although the input of the LLM is text, in most of the architectures, the text (or more generally multi-modal data) input is first decomposed into different tokens and the tokens are processed, but since we consider the LLM as a blackbox, our interpretable models abstract these implementation details.) is given by $\mathcal{D}$, and this dictionary also includes the blank word. The time index is given by $k=1,2,\dots$. A blackbox LLM can be viewed as an input-output block.
@@ III-B Bayesian Utility Maximization Model for LLMA
Following the system prompt, the next input is a context to the LLM. This context could contain external information and examples that are time-dependent and may depend on previous interactions, which is dynamic and can not be put as a part of the system prompt. We consider a context of length $m_{\text{context}}$ and at time $k$ denote it by $\kappa_{k}\in\mathcal{D}^{m_{\text{context}}}$.
@@
We consider an LLM, which is pre-trained on trillions of tokens of text to autoregressively generate the next-token [^6] There are other techniques of generating the token, but they can be accommodated by a suitable augmentation in the mathematical formulation presented here.. For developing a token of length 1, the output of the LLM is given by a conditional probability distribution,
$\phi(c_{k},\kappa_{k},z_{k})=\mathds{P}(\cdot|c_{k},\kappa_{k},z_{k}),$
where $z\in\mathcal{Y}^{\prime}$ is the user prompt, $c_{k}$ is the system prompt and $\kappa_{k}$ is the context. Therefore the function $\phi$,
$\phi:\mathcal{D}^{m_{\text{control}}}\times\mathcal{D}^{m_{\text{context}}}\times\mathcal{D}^{m_{\text{user}}}\to\mathcal{P}(\mathcal{D}),$
outputs a probability distribution over the dictionary $\mathcal{D}$. For generating an output of an LLM with length $>1$, we consider a function $g$ which takes in the function $\phi$ and outputs tokens from the space $\mathcal{D}^{m_{\text{output}}}$, where $m_{\text{output}}$ is the maximum length of the output. The output of the LLM, denoted by $y$ is obtained as $y=g(\phi,c_{k},\kappa_{k},z_{k})$. Therefore, we can represent a black box LLM with the following tuple,
@@
    Current Large Language Models (LLMs) in production lack the ability to explicitly express confidence in their generated outputs. This limitation significantly hinders their deployment in critical domains like finance and healthcare, where system safety and reliability are paramount.
@@
    Furthermore, the absence of explicit confidence scores presents a significant challenge for human-in-the-loop systems. These systems rely on the ability to seamlessly integrate human expertise when an LLM’s confidence in its response falls below a certain threshold. Without this crucial feedback mechanism, it becomes difficult to effectively leverage human oversight. Finally, exclusive reliance on a single LLM for critical tasks is inherently risky. LLMs are vulnerable to adversarial prompting, where carefully crafted inputs can induce them to generate incorrect or misleading outputs. This vulnerability underscores the need for robust mechanisms to assess and mitigate the risks associated with LLM-based systems.
@@
To this end, we propose using the LLM as a map from the high-dimensional space $\mathcal{Y}^{\prime}$ to a low-dimensional space $\mathcal{Y}$ which takes value from whose cardinality $Y\ll\mathcal{D}^{m_{\text{user}}}$. We construct a Bayesian engine composed of two probability distributions, the prior of the state, $\pi\in\mathcal{P}(\mathcal{X})$ and the likelihood of a low-dimensional observation $O:\mathcal{X}\to\mathcal{P}(\mathcal{Y})$ and is a conditional probability distribution. We use $B$ and $O$ to denote the observation matrices in the rest of the paper.
@@
The following example illustrates how LLMs can be used to map the text space to a low-dimensional space. We discuss how this low-dimensional space can be constructed by an analyst or engineer so that the posterior computed by the Bayesian engine is interpretable.
@@ III-B2 Construction of Bayesian Engine
Figure 5: LLMAs can be used to detect and analyze the change in financial indicators (difference of close prices) by parsing financial news to extract 1616 1616 interpretable features in Example 2. We analyze the news articles from $03/2020$ to $08/2020$ corresponding to the AAPL stock. We query the LLM for 1616 1616 binary with different features, including whether the news article indicates optimism about the market and whether there is investor interest in the stock. The interpretable features can be used to analyze the stock (and subsequently for Bayesian inference), as illustrated by the difference in the ratio of the stock prices across days.
@@
For a state $x\in\mathcal{X}$, the LLMA receives a text observation $y^{{}^{\prime}}\in\mathcal{Y}^{\prime}$ which the LLM of the LLMA parses and provides a low-dimensional output $y\in\mathcal{Y}$. The LLMA has a Bayesian engine which has a prior on the state space given by $\pi$ and an observation likelihood $O$. The LLMA uses the low-dimensional output from the LLM to compute the posterior using Bayes’ rule,
$\displaystyle\mathcal{T}(\pi,y)=\mathds{P}_{O,\pi}(x|y)=\frac{O_{y,x}\pi(x)}{\sum_{x}O_{y,x}\pi(x)}.$
@@
$\displaystyle u=\underset{{u\in\mathcal{U}}}{\arg\max}\sum_{x\in\mathcal{X}}r(x,u)\mathds{P}_{O,\pi}(x|y).$
@@
$\displaystyle\text{LLMA}(\mathcal{L},\pi,O,r),$ (2)
where $\mathcal{L}$ is an LLM of the form (1), $\pi$ is the prior over the state, $O$ is the observation likelihood and $r$ is the utility function. We make the following remarks on our interpretable sensor model construction of a LLMA.
@@
We described the operation of Bayesian inference on a single observation when there is a stream of observations $y_{1},\dots,y_{k}$ then the prior is updated after every step with the computed posterior $\pi_{k+1}=\mathcal{T}(\pi_{k},y)$.
@@
The likelihood $O$ can be computed by using the LLM on a set of synthetic or public offline data where we simulate the state and use the text observations to obtain the low-dimensional observations from the LLM.